The market for vanilla ice cream is given by the following information:

𝑄

𝑑 = 800 βˆ’ 30𝑃

𝑣 + 10𝑃

𝑐

𝑄

𝑠 = 250 + 30𝑃

𝑣 βˆ’ 10π‘ƒπ‘š

Where 𝑄

𝑑

is the quantity demanded, 𝑄

𝑠

is quantity supplied, 𝑃

𝑣

is the price of vanilla ice

cream, 𝑃

𝑐

is the price of chocolate ice cream and π‘ƒπ‘š is the price of milk. Let 𝑃

𝑐 = 10 and π‘ƒπ‘š = 5. Calculate the equilibrium price and quantity in the vanilla ice 

cream market. Compare the own-price elasticity of demand and supply at equilibrium. (4


1
Expert's answer
2021-03-22T11:57:16-0400

At equilibrium, supply is equal to demand. So;

Qd=QsQd=Qs

(800βˆ’30Pv+10Pc)=(250+30Pvβˆ’10Pm)(800-30Pv+10Pc)=(250+30Pv-10Pm)

We know that Pc=10Pc=10 and Pm=5Pm=5

hence; 800βˆ’30Pv+10(10)=250+30Pvβˆ’10(5)800-30Pv+10(10)=250+30Pv-10(5)

30Pv+30Pv=800+100+50βˆ’25030Pv+30Pv=800+100+50-250

Pv=12Pv=12 , which the price of Vanilla ice cream at equilibrium.

Therefore, quantity supplied at equilibrium is 250+30(12)βˆ’10(5)250+30(12)-10(5)

=560=560


The quantity demanded at equilibrium will be 800βˆ’30(12)+10(10)800-30(12)+10(10)

=540=540